### v2

# Trajectory-aware Shifted State Space Models for Online Video Super-Resolution

Abstract
Online video super-resolution (VSR) is an important technique for many real-world video processing applications, which aims to restore the current high-resolution video frame based on temporally previous frames. Most of the existing online VSR methods solely employ one neighboring previous frame to achieve temporal alignment, which limits long-range temporal modeling of videos. Recently, state space models (SSMs) have been proposed with linear computational complexity and a global receptive field, which significantly improve computational efficiency and performance. In this context, this paper presents a novel online VSR method based on Trajectory-aware Shifted SSMs (TS-Mamba), leveraging both long-term trajectory modeling and low-complexity Mamba to achieve efficient spatio-temporal information aggregation. Specifically, TS-Mamba first constructs the trajectories within a video to select the most similar tokens from the previous frames. Then, a Trajectory-aware Shifted Mamba Aggregation (TSMA) module consisting of proposed shifted SSMs blocks is employed to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and corresponding shift operations to compensate for scanning losses and strengthen the spatial continuity of Mamba. Additionally, we propose a trajectory-aware loss function to supervise the trajectory generation, ensuring the accuracy of token selection when training our model. Extensive experiments on three widely used VSR test datasets demonstrate that compared with six online VSR benchmark models, our TS-Mamba achieves state-of-the-art performance in most cases and over 22.7% complexity reduction (in MACs). The source code for TS-Mamba will be available at https://github.com/QZ1-boy/TS-Mamba.
1 Introduction
Among various video super-resolution (VSR) application scenarios, online VSR has recently attracted significant interest due to the growing popularity of live video conferencing and live broadcasting applications (Fuoli et al., 2023; Xiao et al., 2023). In online VSR, the current high-resolution (HR) video frame is typically restored using only its low-resolution (LR) counterpart and previous frames. This is constrained by the requirements for low latency and low computational complexity inherent to these online real-time applications.
In a VSR model, temporal alignment or aggregation is a core module employed to compensate for the information from neighboring frames before generating the current HR frame. Advanced temporal alignment or aggregation modules have been recently developed, which are based on deformable convolution networks (DCN) (Wang et al., 2019; Tian et al., 2020), flow-guided deformable alignment models (Chan et al., 2022a; Zhu et al., 2024c), non-local attention mechanisms (Li et al., 2020; Yi et al., 2019), Vision Transformer based spatio-temporal information aggregation (Liu et al., 2022a; Tang et al., 2023) or Diffusion models (Wang et al., 2025; Liu et al., 2025; Zhuang et al., 2025). Although they offer superior VSR performance, these methods are typically associated with high complexity and, therefore, are not ideal for online VSR.
To mitigate these limitations, recent online VSR methods have adopted more efficient temporal alignment modules, such as lightweight optical flow networks (Sajjadi et al., 2018; Xiao et al., 2023), deformable attention mechanisms (Fuoli et al., 2023; Yang et al., 2023), and temporal motion propagation modules (Zhang et al., 2024b). For example, CKBG (Xiao et al., 2023) utilized a lightweight optical flow network to estimate motion between frames and perform motion compensation. DAP (Fuoli et al., 2023) designed a deformable attention pyramid module to dynamically focus on the most salient locations between frames and progressive refine the offsets to improve temporal alignment performance. FDAN (Yang et al., 2023) proposed a flow-guided deformable attention propagation module that introduces the optical flow to guide the offset generation to efficiently exploit the temporal information between frames. Despite their efficiency, these methods predominantly use short-term temporal information based on convolutional neural networks (CNN) — typically from a single previous frame, which restricts their ability to further enhance reconstruction quality. While incorporating long-term temporal alignment can improve performance, it often introduces significant computational overhead, resulting in challenges for real-time or resource-constrained applications. Therefore, it is valuable to develop the efficient long-range models for online VSR applications.
Recently, low-complexity state space models (SSMs) (Gu et al., 2021; Gu & Dao, 2023) have been proposed with linear computational complexity and with relatively large receptive fields, which can potentially improve performance with limited complexity. Inspired by this, we propose a Trajectory-aware Shifted Mamba for online VSR, denoted as TS-Mamba, leveraging long-term trajectory modeling and low-complexity Mamba for achieving the token-level spatio-temporal aggregation. In TS-Mamba, trajectories within a video are first constructed for selecting the most similar tokens from the previous frames. A trajectory-aware shifted Mamba aggregation (TSMA) module is then employed that consists of shifted SSMs blocks to aggregate the selected tokens. The shifted SSMs blocks are designed with specific procedures based on Hilbert scannings and four shift operations to compensate for scanning losses and strengthen the spatial continuity of Mamba. Moreover, we propose a trajectory-aware loss function to supervise the trajectory generation, optimizing the accuracy of token selection when training our model. The proposed TS-Mamba model enables efficient long-term video modeling with significantly reduced computational complexity. The primary contributions are summarized as follows:
TS-Mamba is the first SSMs-based online VSR model, which aggregates long-term spatio-temporal information from previous frames at the token level for restoring current HR frame. This is different from existing online VSR methods which typically use CNN-based temporal alignment to exploit temporal information from a single previous frame.
This is also the first time to introduce video trajectories into Mamba to select the most similar tokens from previous frames and construct the new trajectory-aware shifted Mamba model for efficient token-level spatio-temporal information aggregation.
The novel shifted SSMs blocks are designed based on four different shift operations and Hilbert scannings to effectively compensate for the intra-window and inter-window losses of Hilbert scannings and strengthen the local spatial continuity of Mamba.
The proposed method has been benchmarked on three widely used test datasets and shows superior VSR performance with more than 22.7% computational complexity reduction in terms of MACs over five state-of-the-art (SOTA) online VSR methods (as shown in Figure 1).
2 Related Work
2.1 Video Super-Resolution
Video super-resolution (VSR) is a fundamental low-level vision task that aims to restore an HR video from its LR counterpart. Existing VSR methods are typically learning-based, utilizing various deep neural networks (Teed & Deng, 2020; Zhu et al., 2019; Arnab et al., 2021; Ho et al., 2022). For example, optical flow-based methods (Chan et al., 2021; Liu et al., 2022b) explore the temporal motion between frames to align them; deformable convolution networks (DCN)-based methods (Tian et al., 2020; Wang et al., 2019; Dong et al., 2023) learn the motion offsets between frames for feature alignment. Moreover, flow-guided deformable-based methods (Chan et al., 2022a; Zhu et al., 2024c) combine optical flow and DCN to achieve better feature alignment. Non-local attention-based methods (Li et al., 2020; Yi et al., 2019) aggregate global information for feature aggregation. Vision Transformer-based methods (Liu et al., 2022a; Tang et al., 2023; Lin et al., 2022) aggregate long-term spatio-temporal information in video to restore SR frames. Diffusion-based methods (Wang et al., 2025; Liu et al., 2025; Zhuang et al., 2025) build the long-range modeling using designed diffusion models for information aggregation. However, these methods are still associated with high complexity and are therefore not best suited for real-time online VSR applications. Recently, some Mamba-based works (Xiao & Wang, 2025; Tran et al., 2025) were proposed that use low-complexity Mamba with global receptive filed to improve VSR performance. Although Mamba-based methods achieve some reduction in model complexity, their high-overhead repeated scannings hinders efficient implementation of real-time online VSR.
2.2 Online Video Super-Resolution
Due to the specific requirements of online applications, online VSR methods are expected to be lightweight and have low latency. Therefore, most existing online VSR methods have been proposed (Sajjadi et al., 2018; Cao et al., 2021; Fuoli et al., 2023; Xiao et al., 2023; Jiang et al., 2025a) with efficient feature alignment modules. For example, EGVSR (Cao et al., 2021) and CKBG (Xiao et al., 2023) utilized lightweight optical flow networks to estimate motion between frames and perform motion compensation. KSNet (Jin et al., 2023) proposed a kernel-split manner to reparameterize convolutional kernels on the high-value channel, enabling representation of dynamic information and reducing complexity along the channel dimension. TMP (Zhang et al., 2024b) employs an efficient temporal motion propagation method that leverages motion field continuity to achieve fast feature alignment. DAP (Fuoli et al., 2023) designed a deformable attention pyramid module to dynamically focus on the most salient locations between frames and progressive refine the offsets to achieve temporal alignment improvement. FDAN (Yang et al., 2023) proposed a flow-guided deformable attention propagation module that introduces the optical flow to guide the offset generation to efficiently exploit the temporal information between frames. It is noted that, however, these online VSR methods are only based on one previous frame in feature alignment due to the complexity limitation, which hinders further improvement of VSR performance. Different from existing deformable-based methods (Fuoli et al., 2023; Yang et al., 2023) that based on short-term spatio-temporal aggregation, in this work, our TS-Mamba introduces of trajectories and designs shifted SSMs blocks, enabling it improve the ability for long-range spatio-temporal information aggregation.
2.3 State Space Models
State space models (Gu et al., 2021; Gu & Dao, 2023), e.g., Mamba, have been widely employed in vision tasks (Liu et al., 2024; Zhu et al., 2024a) due to their linear computational complexity and ability to model global dependencies. Recently, some Mamba-based methods are proposed for image/video super-resolution. For example, MambaIR (Guo et al., 2024), and MambaIRv2 (Guo et al., 2025) used Mamba to achieve the global receptive field. TAMambaIR (Peng et al., 2025) introduced a texture-aware state space model to focus on textures regions for improving SR quality. VSRM (Tran et al., 2025) proposed the spatial-to-temporal Mamba and the temporal-to-spatial Mamba to ability of spatio-temporal aggregation. MamEVSR (Xiao & Wang, 2025) proposed a interleaved Mamba and a cross modality Mamba to interleave tokens and further leverage spatio-temporal information to capture finer details. Typically, Mamba converts 2D images into 1D tokens through scanning (Qiao et al., 2024; Shi et al., 2025), resulting in spatial continuity loss inherent to images. To enhance the ability of Mamba, some advanced scanning techniques have emerged to address this issue, such as bidirectional scanning (Hu et al., 2024; Shi et al., 2025), cross scanning (Liu et al., 2024), continuous 2D scanning (Yang et al., 2024a), and local scanning (Huang et al., 2024). To the best of our knowledge, the use of Mamba has not yet been investigated for the online video super-resolution task. Unlike existing Mamba-based works (Xiao & Wang, 2025; Tran et al., 2025) that neglect the local spatial continuity of Mamba, we introduce sophisticated shift operations for Hilbert scannings to enhance the ability of Mamba to maintain local spatial continuity, improving the online VSR performance with high efficiency.
3 Methodology
In online video super-resolution, when reconstructing the frame in a low-resolution video, we denote the current LR frame as and temporally previous LR frames as . The proposed trajectory-aware shifted state space models, TS-Mamba, are illustrated in Figure 2. Here, all these LR video frames are first fed into the token and trajectory generation module to extract the current frame token and the tokens of previous LR frames :
where consists of a convolution layer and residual blocks to generate features and tokens from video frames, is the token number, and is the temporal window size. Based on the generated tokens , the trajectories of frame can be formulated as a set of trajectories,
where , and and represent the height and width of the feature (for LR frame), respectively. Each trajectory contains a sequence of coordinates , and the end point of trajectory is associated with the coordinate of token .
We then select the most similar tokens along the trajectories and feed them into the proposed trajectory-aware shifted Mamba aggregation (TSMA) module alongside token to achieve spatio-temporal information aggregation:
Finally, the aggregated feature and current LR frame are fed into the reconstruction network and the upsampling network, respectively, to produce the super-resolved frame :
in which consists of two convolution layers, residual blocks, and a pixelshuffle layer. here represents the bicubic upsampling operation.
3.1 Token Selection
In order to select the most similar tokens along trajectories, we first reformulate tokens , associated with trajectories . Based on the formulation of the trajectories in Equation 3, tokens , and can be formulated as:
We compute cosine similarity between the token , and tokens to select the most similar tokens along trajectories. The indices of the selected tokens and the selected tokens can be formulated as:
Thus, the process of the TS-Mamba network is described as:
3.2 Trajectory-aware Shifted Mamba Aggregation
Mamba networks are typically used to convert 2D images into 1D tokens via scanning, resulting in spatial continuity losses inherent to the images. Existing works (Zhang et al., 2024a; Xiao & Wang, 2025) do not analyze the degree of discontinuous regions but instead repeatedly use multiple scannings, making these methods hard to maintain the spatial continuity of the image and instead lead to greater complexity.
[…]
3.3 Discontinuity for Hilbert Scanning
To evaluate the spatial discontinuity of Hilbert scannings in local windows, we define the discontinuity degree as follows. If the four adjacent areas are successively scanned, the region consisting of these four scanned areas is considered as continuous ( = 0); otherwise, the discontinuity degree equals the number of areas that are not successively scanned. For a region consisting of four adjacent areas, the range of discontinuity degree is . This is illustrated in Figure 3 (a), where four typical Hilbert scannings, i.e., , Scan-2, Scan-3, Scan-4}, are shown on a 44 grid. Here the region with = 1 is marked by a green circle and the region with = 2 is marked by a red circle.
Moreover, we extend the general case to that based on the 88 grid to further discuss the discontinuity degrees. An 88 region is partitioned into four 44 local regions and we illustrate the discontinuity degree within and between local windows under Scan-1 in Figure 3 (b). It can be observed that both intra-window discontinuity and inter-window discontinuity exist. In particular, due to the nature of Hilbert scanning, the central region between windows is widely spaced (the inter-window discontinuity), resulting in inter-level gaps. Here, the discontinuity degree equals 3 - we mark this region with a gray circle in Figure 3 (b).
3.4 Shifted SSMs Block
To eliminate the discontinuity of Hilbert scannings, we propose the “Scan-Shift-Scan” manner that combines window shifting with specific Hilbert scannings to strengthen the continuity of SSMs. The shifting can be defined based on the shift direction and shift position, e.g., Up 1 position ((1)), Up Left 1 position ((1)) and Down Right 2 position ((2)). Our “Scan-Shift-Scan” manner is designed based on the four Hilbert scannings (shown in Figure 3 (a)) and these window shifting processes. As shown in Figure 3 (b), we illustrate the procedure of Scan-1Scan-3 as an example. The local windows are first partitioned by shift operation and then cyclic fed as the shifted windows. It can be inferred that the second scanning (Scan-3) on the shifted window can eliminate the discontinuity of first scanning (Scan-1). To simplify the description of procedure, we define the procedure as:
where the first and the second scannings are denoted as , . Shift operations are denoted as . Therefore, procedure Scan-1Scan-3 can be formulated as .
To evaluate the discontinuity elimination, we set three symbols and define an elimination value to mark and calculate the elimination. Specifically, we use “\”, “\”, and “\” on the circle for representation that eliminates 1, 2, and 3 discontinuity degrees, respectively, in Figure 3 (b)-(c). The elimination value is calculated by summing the eliminated discontinuity degrees that consist of intra-window discontinuity elimination and inter-window discontinuity elimination, i.e., . We have investigated possible procedures and illustrated the representative shifted windows generated by six shift operations, i.e., (1), (2), (3), and (1), (2), (3), under the first scanning (Scan-1) and second scanning (Scan-3) in Figure 3 (c). It can be inferred from Figure 3 (b)-(c) that procedure achieves the best elimination (=18), and the best intra-window discontinuity elimination (=18) but doesn’t eliminate inter-window discontinuity (=0). We infer that the other three procedures also achieve the best elimination: , , . Moreover, procedures , and have the best inter-window discontinuity elimination (=6) but worse than procedure for intra-window discontinuity elimination (=8). The more procedures and details are provided in the supplementary.
[…]
3.5 Selective Scanning along Temporal Dimension
To achieve temporal token aggregation, we implement spatial Hilbert-based selective scanning along the temporal dimension, i.e., SS3D. As shown in Figure 2, we showcase the SS3D processing with Scan-1. The current token and selected tokens are scanned to convert spatio-temporal neighboring pixels into a 1D token sequence. Each token sequence undergoes selective scanning based on the local windows. This process interweaves selected tokens with current tokens, enabling information to interact across spatial and temporal dimensions to capture long-term spatio-temporal characteristics. By scanning spatio-temporally adjacent pixels, SS3D preserves local spatial information and progressively captures global temporal patterns.
3.6 Loss Function
We adopt Charbonnier loss (Lai et al., 2018) as the spatial loss function to supervise the SR frame generation:
in which is the HR frame and the is set to .
To supervise the trajectory generation for ensuring the accuracy of token selection, we first employ the formulation of trajectories of LR video in Equation 3 to generate trajectories of HR video:
Based on this, we propose our trajectory-aware loss function:
where is the downsampling with scale factor that subsamples every coordinate to LR size.
Overall, the total loss is:
in which the hyperparameter is set to 0.1.
4 Experiments
4.1 Experimental Settings
Following previous online VSR research (Jin et al., 2023; Zhang et al., 2024b), we use REDS (Nah et al., 2019), and Vimeo-90K (Xue et al., 2019) as training datasets. REDS4 is used for evaluating the models trained on the REDS dataset, while Vimeo-90K-T and Vid4 (Liu & Sun, 2013) are utilized for benchmarking the models trained on the Vimeo-90K dataset. Two degradations, BI (bicubic) and blur degradation (BD), are used to perform downsampling and the downsampling factor is set to = 4. For BI downsampling, the HR frame is downsampled by a bicubic filter. For BD downsampling, the HR frame is first blurred by a Gaussian filter with standard deviation = 1.6, and then the blurred frame is subsampled for every pixels to generate the LR frame. PSNR and SSIM are adopted as performance evaluation metrics. Runtime (Run.), FPS (frames per second), MACs, and parameters (Params.) are computed on an LR frame of size 180320 to evaluate model complexity and speed.
In the experiments, the numbers of residual blocks and are set to 2 and 13, respectively. The token size is 44 and the window size is 88. The selected token number is set as 3. Random flips, rotations, and temporal inversion operations are performed for data augmentation. Adam optimizer (Kingma, 2014), and Cosine Annealing scheme (Loshchilov & Hutter, 2016) are used during network training. The HR patch size is 256256 and the batch size is 8. The total number of iterations is 600K. The proposed method is implemented on the PyTorch platform with two NVIDIA GeForce RTX 3090 GPUs. Following (Liu et al., 2022a), a lightweight optical flow network (Kong et al., 2021) is adopted to update trajectories. The temporal window size is set as 15 based on (Zhang et al., 2024b) when training on REDS (Nah et al., 2019). For the Vimeo-90K (Xue et al., 2019) dataset, the original sequence is temporally flipped to obtain a 14-frame sequence.
We compare our approach with five SOTA online VSR methods, including RRN (Isobe et al., 2020), DAP-128 (Fuoli et al., 2023), FDAN (Yang et al., 2023), KSNet (Jin et al., 2023), and TMP (Zhang et al., 2024b), and eight bidirectional propagation VSR methods, BasicVSR (Chan et al., 2021), IconVSR (Chan et al., 2021), BasicVSR++ (Chan et al., 2022a), SSL (Xia et al., 2023), DFVSR (Dong et al., 2023), MIA-VSR (Zhou et al., 2024b), IART (Xu et al., 2024) and VSRM (Tran et al., 2025)). Additionally, we implemented another methods, i.e.,“BasicVSR++*” and “VSRM*”, by removing the backward propagation branch of VSR models, i.e., BasicVSR++ and VSRM, and reducing its model size for online VSR application. We use “P”,“F” and “N” to represent those with the previous support (supp.) frames, future support frames and no support frames.
| Category | Methods | Supp. Frame | R-T. | Run. (ms) | FPS (1/s) | MACs (G) | Params. (M) | BI degradation | BD degradation | ||
| REDS4(RGB) (PSNR/SSIM) | Vid4(Y) (PSNR/SSIM) | Vimeo-90K-T(Y) (PSNR/SSIM) | Vid4(Y) (PSNR/SSIM) | ||||||||
| Bidirectional | BasicVSR | P+F | ✗ | 63 | 15.9 | 397 | 6.3 | 31.42/0.8909 | 27.24/0.8251 | 37.53/0.9498 | 27.96/0.8553 |
| IconVSR | P+F | ✗ | 70 | 14.3 | 452 | 8.7 | 31.67/0.8948 | 27.39/0.8279 | 37.84/0.9524 | 28.04/0.8570 | |
| BasicVSR++ | P+F | ✗ | 77 | 13.0 | 418 | 7.3 | 32.39/0.9069 | 27.79/0.8400 | 38.21/0.9550 | 29.04/0.8753 | |
| SSL-bi | P+F | ✗ | 24 | 41.7 | 92 | 1.0 | 31.06/0.8933 | 27.15/0.8208 | 37.06/0.9458 | 27.56/0.8431 | |
| DFVSR | P+F | ✗ | - | - | - | 7.1 | 32.76/0.9081 | 27.92/0.8427 | 38.51/0.9571 | 29.56/0.8983 | |
| MIA-VSR | P+F | ✗ | 318 | 3.1 | 3220 | 16.5 | 32.78/0.9220 | 28.20/0.8507 | - | - | |
| IART | P+F | ✗ | 180 | 5.6 | 5020 | 13.4 | 32.90/0.9138 | 28.26/0.8517 | 38.62/0.9579 | 29.68/0.8884 | |
| VSRM | P+F | ✗ | 223 | 4.5 | 2174 | 17.1 | 33.11/0.9162 | 28.44/0.8552 | - | - | |
| Online | Bicubic | N | ✓ | - | - | - | - | 26.14/0.7292 | 23.78/0.6347 | 31.30/0.8687 | 21.80/0.5246 |
| RRN | P | ✓ | 34 | 29.4 | 193 | 3.4 | 28.82/0.8234 | 25.85/0.7660 | 36.69/0.9432 | 27.69/0.8488 | |
| BasicVSR++* | P | ✓ | 40 | 25.0 | 146 | 3.0 | 30.44/0.8686 | 27.06/0.8173 | 37.11/0.9464 | 27.49/0.8426 | |
| DAP-128 | P | ✓ | 38 | 26.3 | 165 | - | 30.59/0.8703 | - | 37.29/0.9476 | - | |
| FDAN | P | ✓ | 34 | 29.4 | 146 | 3.9 | 30.71/0.8723 | 27.14/0.8206† | 37.36/0.9483† | 27.76/0.8471 | |
| KSNet | P | ✓ | 31 | 32.3 | 145 | 3.0 | 30.69/0.8724 | 27.14/0.8208 | 37.34/0.9490 | 27.63/0.8444† | |
| TMP | P | ✓ | 25 | 40.1 | 176 | 3.1 | 30.67/0.8710 | 27.10/0.8167 | 37.33/0.9481 | 27.61/0.8428 | |
| VSRM* | P | ✓ | 31 | 32.7 | 136 | 3.1 | 30.64/0.8701 | 27.10/0.8163 | 37.28/0.9477 | 27.57/0.8423 | |
| TS-Mamba | P | ✓ | 29 | 33.5 | 112 | 3.0 | 30.73/0.8727 | 27.17/0.8209 | 37.36/0.9482 | 27.70/0.8473 | |
4.2 Overall Performance
As shown in Table 1, the quantitative results demonstrate the superior performance of the proposed method over other online VSR models in terms of PSNR and SSIM. We also supplement the results of FDAN and KSNet models on Vid4 and Vimeo-90K-T datasets based on their released pre-trained models and source codes for a comprehensive comparison. These results are reported in Table 1 with “†”. Figure 4 presents qualitative comparisons, from which we can observe that our method shows better visual quality than other online VSR methods for both BI and BD degradations.
REDS4_011_011 (BI)
GT
BasicVSR++*
FDAN
KSNet
TMP
TS-Mamba
Vid4_City_009 (BI)
GT
BasicVSR++*
FDAN
KSNet
TMP
TS-Mamba
Vimeo90K-T_001_0837(BD)
GT
BasicVSR++*
FDAN
KSNet
TMP
TS-Mamba
Vid4_Foliage_025 (BD)
GT
BasicVSR++*
FDAN
KSNet
TMP
TS-Mamba
Following (Fuoli et al., 2023; Zhang et al., 2024b), VSR methods that can process 720p (1280720) videos in at least 24 in terms of FPS are recognized as real-time (R-T.) methods (Fuoli et al., 2023), and we have labelled all the tested methods in Table 1 according to their runtime. It is noted that our TS-Mamba model achieves the second fastest inference speed among all online VSR methods. TMP is the one with the fastest runtime as it was implemented with the CUDA accelerator (high MACs but low runtime) while TS-Mamba is not. Moreover, TS-Mamba also offers a significant reduction in terms of MACs (about 36.3%) and a marginal reduction in parameter numbers compared to TMP, as shown in Figure 1.
4.3 Ablation Study
To further verify the effectiveness of our contributions, we have conducted ablation studies on the REDS4 dataset with BI degradation.
| Models | PSNR(dB) / SSIM | Params.(M) | Run.(ms) | MACs(G) |
| (v1.1) w/o Trajectory | 30.45 / 0.8678 | 1.7 | 20 | 84 |
| (v1.2) w/o | 30.70 / 0.8721 | 3.0 | 29 | 112 |
| (v1.3) w/o IntraWCB | 30.58 / 0.8702 | 2.8 | 25 | 97 |
| (v1.4) w/o InterWCB | 30.61 / 0.8706 | 2.8 | 25 | 97 |
| (v1.5) w/o IntraWCB+InterWCB | 30.52 / 0.8689 | 2.4 | 21 | 85 |
| (v1.6) w/o | 30.65 / 0.8710 | 3.0 | 27 | 112 |
| (v1.7) w/o | 30.67 / 0.8714 | 3.0 | 27 | 112 |
| (v1.8) w/o (v1.6) + (v1.7) | 30.61 / 0.8702 | 3.0 | 25 | 111 |
| TS-Mamba (ours) | 30.73 / 0.8727 | 3.0 | 29 | 112 |
| GT | w/o Trajectory | w/o | w/o IntraWCB | w/o InterWCB | |
| REDS4_000_045 (BI) | w/o IntraWCB + InterWCB | w/o | w/o | w/o (v1.6)+(v1.7) | TS-Mamba (ours) |
We first confirmed the contribution of two trajectory-aware designs, i.e., trajectory generation and trajectory-aware loss, by creating the following variants. (v1.1) w/o Trajectory - and Token Selection module were removed from TS-Mamba; (v1.2) w/o - the trajectory-aware loss function was removed when training the TS-Mamba model. We further verified our proposed TSMA module in terms of compensation branches and shift operations, by obtaining (v1.3) w/o IntraWCB and (v1.4) w/o InterWCB - IntraWCB and InterWCB were removed from the TSMA module, respectively; (v1.5) w/o IntraWCB + InterWCB - both IntraWCB and InterWCB were removed from the TSMA module. We also tested the adopted shift operations in compensation branches, and implemented (v1.6) w/o - the shift operations were removed in IntraWCB, (v1.7) w/o - the shift operations were removed in InterWCB and (v1.8) w/o (v1.6)+(v1.7) - all the shift operations were removed in TSMA module. As shown in Table 2, the performance of all these variants is evidently lower than that of the full TS-Mamba, which fully confirms the effectiveness of each key component in our design. Moreover, we further provide the visual results of these variants on REDS4 dataset in Figure 5. It is found that the visual results demonstrates the contributions of our designs, particularly in realistic textures and fine details of the car.
| PSNR(dB)/SSIM | Params. (M) | Run. (ms) | MACs (G) | |
| 1 | 30.64/0.8712 | 2.8 | 25 | 96 |
| 2 | 30.68/0.8720 | 2.9 | 27 | 104 |
| 3 | 30.73/0.8727 | 3.0 | 29 | 112 |
| 4 | 30.74/0.8727 | 3.1 | 31 | 120 |
REDS4_015_049
GT
FDAN
TMP
VSRM*
TS-Mamba
To confirm the value of the token number in our TS-Mamba, we tested different values with our TS-Mamba, and presented the results in Table 3. It is noted that as increases, the VSR performance improves, but with higher model complexity. When = 4, it is difficult to obviously improve VSR performance. To trade off between complexity and performance, we set = 3 in this work.
5 Limitations
We investigate our results and find out the failure cases. A failure case when highly dynamic rotation occurs is visualized in Figure 6. The generated trajectories and compensated manner of our TS-Mamba are inaccurate enough when dynamic rotation occurs in car tire and rotation information cannot be reconstructed, thus limiting the performance of our method. Due to the high difficulty of modeling rotation, other online VSR methods also fail to obtain complete rotation information.
6 Conclusion
In this paper, we proposed a Trajectory-aware Shifted SSMs (TS-Mamba) for online VSR, leveraging long-term trajectory modeling and low-complexity Mamba to achieve efficient spatio-temporal information aggregation. In TS-Mamba, trajectories in a video are first constructed to select the most similar tokens from the previous frames. A trajectory-aware shifted Mamba aggregation module is then employed, which consists of shifted SSMs blocks to aggregate the selected tokens. The shifted SSMs blocks are designed based on Hilbert scannings and shift operations to compensate for the scanning losses and strengthen the spatial continuity of Mamba. Extensive experiments on three widely used VSR benchmarks have demonstrated the effectiveness and efficiency of our method.